Albright Properties Inc. (API) has three divisions:

Division Beta Proportion of Firm’s Assets
Property management 1.0 30%
Land resources 1.5 40
Financial services 30

The leveraged beta for API is 1.1. API has a consolidated capital structure consisting of 50 percent debt and 50 percent equity. The Financial Services Division’s capital structure is 80 percent debt and 20 percent equity. API is planning to finance new projects in that division with a capital structure of 90 percent debt and 10 percent equity. The risk-free rate is 5 percent, and the market risk premium is 8.9 percent. The pretax cost of debt to API is 20 percent. The tax rate is 40 percent. What discount rate should API apply to the cash flows from “new” projects in the Financial Services Division? Round your answer to one decimal place.

%
Answer #1

The beta of the API is 1.1 1*0.30 + 1.5 * 0.40 + Beta *0.30 = 1.1 beta * 0.30 = (1.1-(1*0.3 +1.5*0.4)) 0.20 beta = (0.2/0.3)